OpenStudy (anonymous):

Please help!!! Urgent!! A department store is having a clearance sale. The price on a television is reduced by 38%. That sale price is then reduced by another 38%. If x is the televisions original price, the sale price can be represented by (x-0.38x)-0.38(x-0.38x). With these two reductions, at what percentage of the original price is the television being sold? Use the factored, simplified form of the expression to answer the question.

OpenStudy (anonymous):

@triciaal

OpenStudy (anonymous):

@zepdrix

OpenStudy (anonymous):

@ganeshie8

OpenStudy (anonymous):

@Compassionate

OpenStudy (anonymous):

@texaschic101

OpenStudy (triciaal):

The price on a television is reduced by 38%. full price x = 100% sale price = 62% *x

OpenStudy (anonymous):

Soo 100 would go into x

OpenStudy (triciaal):

That sale price is then reduced by another 38%. new price = 62% of 62% * x

OpenStudy (anonymous):

I would subtract 100 to get 62?

OpenStudy (triciaal):

Use the factored, simplified form of the expression to answer the question. so use a different approach should get the same answer.

OpenStudy (anonymous):

38.44%is the answer

OpenStudy (anonymous):

May you please help me with another question?

OpenStudy (triciaal):

(x-0.38x)-0.38(x-0.38x) = 62/100

OpenStudy (triciaal):

how did you get that and was it correct?

OpenStudy (anonymous):

Yes that answer was correct Solve the quadriatic equation 7/(x-5) + x/(x+5)= 44/(x+5)(x-5)

OpenStudy (triciaal):